# Towards Immersive Mixed Reality Street Play

### v2

@@ 1. Introduction
Major technology companies such as Apple and Meta are rapidly advancing see-through Mixed Reality Head-Mounted Display (MRHMD) technology: Apple’s Vision Pro (Apple Inc., 2023) offers video see-through mixed reality in a headset form factor, while Meta’s Orion (Meta Platforms, Inc., 2024) boasts optical see-through capabilities in an eyeglasses‑style form factor. Meta CEO Mark Zuckerberg has even predicted that “smart glasses will replace phones by 2030” (Birch, 2024). Some scholars forecast that augmented reality could become virtually indistinguishable from physical reality in certain contexts (Nijholt, 2023; Laato et al., 2021; Rauschnabel, 2021), underscoring the importance of understanding user experiences and social dynamics surrounding these emerging devices. As Itoh et al. (Itoh et al., 2021) demonstrate, passthrough technologies—whether video see-through or optical see-through—enable users to maintain awareness of their real-world surroundings while moving freely, extending potential usage from controlled private settings to spontaneous public scenarios. This shift raises pressing questions about how ubiquitous MRHMD adoption could transform everyday public life in the near future.
@@ 2.1. Technology-mediated Street Play
Informed by such studies, HCI has developed conceptual frameworks for thinking about designing for interaction in public, including research in the wild (Chamberlain et al., 2012), performance-led research (Benford et al., 2013), and design considerations of bystanders and spectators in public interactions (Reeves et al., 2005; Reeves, 2011). These approaches have been instrumental in understanding and shaping user experiences in outdoor settings. Meanwhile, HCI research has documented numerous social phenomena that arise when people interact with technology in public contexts, such as public displays (O’Hara et al., 2008) and public installations (Balestrini et al., 2016). Among these are the honeypot effect (Wouters et al., 2016), where an initial user attracts onlookers and encourages them to engage; showmanship and performance, where users become aware of an audience and amplify their actions  (Benford and Giannachi, 2011; Peltonen et al., 2008; Reeves et al., 2005); social mirroring, in which newcomers imitate observed behaviors (Peltonen et al., 2008). Researchers have also noted social discomfort (Benford et al., 2012), social facilitation (improved performance in front of others) (Peltonen et al., 2008; Miller et al., 2019), social awkwardness (Deterding, 2018; Huggard et al., 2013; Press and Erel, 2023), and inhibition (anxiety or reluctance to participate publicly) (Miller et al., 2019).
@@ 2.2. Technology-supported Co-located Bodily Play
In recent years, HCI research has undergone a “somatic turn” (Loke and Schiphorst, 2018). This shift places the human body at the center of interactive technology design, moving beyond traditional user-centered approaches (Dix, 2017) to focus on the body as the primary site of interaction and experience (Höök, 2018). Merleau-Ponty argues that the human body is fundamental to perception and action (Merleau-Ponty, 2008), with proprioception and kinesthesia serving as core movement-related senses (Zhou et al., 2021). Homewood et al. (Homewood et al., 2021) outline three evolving conceptions of the body in HCI: the individual body (embodied interaction), multiple bodies (plurality of experiences), and more-than-human bodies (entangled assemblages). These concepts, particularly entangled assemblages (Forlano, 2017), intercorporeality (Meyer et al., 2017; Stepanova and Reicke, 2024; Patibanda et al., 2024), and post-phenomenology (Wakkary et al., 2018), provide theoretical frameworks for exploring technology-supported co-located bodily play.
@@ 2.4. Mixed Reality Concerns in Public Spaces
In recent years, MRHMDs with passthrough functionality (Itoh et al., 2021) have sparked a notable shift from private, controlled settings to more public, spontaneous environments. These devices, including video and optical see-through technologies like Apple Vision Pro, Magic Leap, Microsoft HoloLens, and Meta Orion, enable users to view their physical surroundings while interacting with virtual elements (Radu et al., 2021; Bailenson et al., 2024). The rise in MRHMD usage in public settings raises social concerns for both wearers and non-wearers (Gugenheimer et al., 2017; Sapkota et al., 2021). The social acceptance (Koelle et al., 2020) of MRHMD usage in public spaces tends to decline significantly (Koutromanos and Kazakou, 2023; Profita et al., 2016; Sun and Schöning, 2025; Schwind et al., 2018; Eghbali et al., 2019; Gugenheimer et al., 2019, 2022; O’Hagan et al., 2023; Lebeck et al., 2018) due to three main factors:
@@
MRHMDs grant wearers capabilities—such as data overlays or video recording—that non-wearers cannot see or fully understand. This imbalance can negatively affect social experiences, privacy, and safety (Chung et al., 2023). Denning et al. (Denning et al., 2014) examined bystander privacy in AR, while O’Hagan et al. (O’Hagan et al., 2021, 2022) studied bystanders’ privacy concerns and awareness regarding AR glasses. A notable example of these social tensions is Google Glass, which faced criticism for prioritizing user capabilities over broader social impacts. This led to the term “Glasshole”, illustrating bystanders’ perception of antisocial behavior by Google Glass wearers (Due, 2015).
@@
When MRHMDs are used “on the go” in public spaces (Schwind et al., 2018), such as urban streets (Guo et al., 2024), public libraries (Kaeder et al., 2024), public transit (Bajorunaite et al., 2025, 2023; Medeiros et al., 2022; Bajorunaite et al., 2024; Medeiros et al., 2024a), and airplanes (Ng et al., 2021; Williamson et al., 2019), social acceptance, trust, security, and privacy become even more pressing (Katins et al., 2024; Millard et al., 2024; Roesner et al., 2014; Zhu et al., 2024b; Abraham et al., 2022). Interactions and movement in public traffic environments can pose additional risks, especially when MRHMD wearers navigate between physical and virtual elements. This interaction style intensifies the power imbalance and increases the potential for accidents or social friction with bystanders.
@@ 2.6. Methods for Exploring Future Mixed Reality
Traditional design futuring (Fry, 2018; Kozubaev et al., 2020; Celik and Kaya, 2025; Lindley and Coulton, 2015) practices, including speculative design (Auger, 2013; Dunne, 2013; Barendregt and Vaage, 2021), design fiction (Sterling, 2009; Blythe, 2014), experiential futures (Candy and Dunagan, 2017), critical design (Bardzell et al., 2012), material speculation (Wakkary et al., 2015), typically employ fictional or low‑fidelity artifacts (Zhu et al., 2024a; Gaver, 2012), such as diegetic prototypes (Kirby, 2010), mock‑ups (Blythe et al., 2016), scripted scenarios (Candy and Dunagan, 2017), games (Coulton et al., 2016; Brown et al., 2016), or concept videos (Briggs et al., 2012; Odom et al., 2024; Sturdee et al., 2016) to provoke reflection, to imagine alternative worlds, and to foreground the discursive power. Such formats, however, often leave audiences on the far side of what Candy calls an “experiential gulf” (Candy, 2010): viewers discuss a hypothetical future rather than live it, producing debate rather than situated behavioral data (Ringfort-Felner et al., 2025). Elsden et al. (Elsden et al., 2017) further caution that lab‑based user enactments (Odom et al., 2012) can break the suspension of disbelief. To address this gap and facilitate experiential participation, Speculative Enactments (Elsden et al., 2017) stage scenarios as lived experiences that foreground real‑world consequences. Building on this, Immersive Speculative Enactments (Simeone et al., 2022) leverage virtual reality technologies to place participants in open‑ended situations that invite authentic action rather than following fixed linear narratives, as is common in Immersive Design Fiction (McVeigh-Schultz et al., 2018). Cools et al. (Cools et al., 2024) offer an Immersive Speculative Enactment example in which participants experience fictional MR contact lens scenarios through a multi-user VR simulation, generating experiential insights into the technology’s potential impact on daily life.
@@ 3.3. System Implementation
We connect all co-located devices using the InstantCopresence toolkit (Hu et al., 2023a). The toolkit enables spontaneous mixed reality sessions when users scan a QR code on the host device. QR code scanning enables device co-location by calculating the relative position of client devices to the host’s QR code, allowing multiple devices to sync their coordinate systems based on their positions relative to the host origin. The network layer uses Multipeer Connectivity—Apple’s peer-to-peer wireless Wi-Fi and Bluetooth technology that powers AirDrop—to provide local networking with latency under 20 ms. This enables street play anywhere, even in rural areas without cellular or Wi-Fi coverage. An auto-sync system keeps virtual object data consistent across devices, ensuring all participants see the same virtual elements. Spectators can view the augmented content from a third-person perspective using mobile devices like iPads (see Fig. 4).
@@ 4.1.2. Procedure
    Conclude gameplay: Participants could play for their desired duration. However, as the MOFA games involved vigorous physical activity, players typically concluded after 3–5 minutes due to physical fatigue from bodily exertion. When bystanders wished to join, researchers rotated them into the session. This fluid grouping mimicked real-world street play, where people can spontaneously join or leave a game.
@@ 4.1.4. Safety and Ethical Considerations
We avoided any hazardous locations. Although privacy is not expected in public spaces, we followed strict ethical guidelines during data collection. The focus of our capture was the play, and we avoided recording people unnecessarily. Participants or bystanders could opt out before interviews. No minors were interviewed; a few minors who tried the game did so under parental supervision, but we did not include their feedback in our dataset beyond general observations.
@@ 4.2. Desktop Research of Online Forum Reviews
In addition to on-site data, we wanted to capture broader public perceptions of IMRSP—including those of people who had not played the game but saw it online. To do this, we conducted a desktop research analysis of online discussions about our game probe. We posted a short gameplay video of MOFA and a description of the concept on a popular tech-enthusiast forum, inviting feedback and discussion. Over several weeks, this thread garnered considerable attention and commentary. We collected all responses from this forum thread, resulting in 14,919 words of raw text across numerous comments. Fig. 7 shows some snippets of the online forum threads. We then filtered this dataset to focus on relevant content, specifically, comments that addressed the social implications, user experience, or concerns about MR street play (as opposed to off-topic remarks about technology or economics). After filtering out unrelated discussion, we had 1,325 words of substantive comments from 31 unique online respondents (L1–L31). These comments ranged from short quips to multi-sentence opinions (5 to 204 words each).
The online respondents had all viewed the same gameplay footage that our real-world participants had experienced (via the posted video). Their perspectives are thus those of observers envisioning the experience, rather than firsthand players. Not being physically present or involved with researchers, their feedback was not influenced by any demand characteristics or desire to please us. This provided us with more candid feedback and skepticism than we might have received from the public onsite. Some commenters were enthusiastic about the idea, while others were critical or concerned. We treated this as an additional qualitative dataset that complements the interview data: on-site participants provide insight into actual experience and immediate personal feelings, whereas online responses reveal more reflective or hypothetical viewpoints, including critiques from a social acceptance standpoint.
@@ 5.2.1. Social awkwardness and perceived impropriety (13/63)
Thirteen respondents (13/63), all from the online forum data collection, expressed feelings of awkwardness as their main impression of the game, with some feeling offended by the prospect of engaging in immersive MR street play. These comments questioned the social acceptance level of wearing the device in public, particularly regarding the potential impact on non-participants. For example, “I’m not sure if I’d like to walk around the streets wearing them though—it would likely be incredibly annoying to other people” (L1). L19 criticized the play as “disregard for culture and obvious social media reactions.” One comment expressed empathy for passersby who might be unintentionally captured in the middle of others’ game space, noting that they “don’t look too pleased” (L26). The impropriety of engaging in the game in certain contexts was also highlighted, with L10 commenting, “It seems pretty impolite to be LARPing in a busy train station”. The remaining comments expressed more straightforward emotions towards this phenomenon, accusing the MR street players of being “a nuisance to everyone” (L27), and finding it to be “complete idiots to play in public places” (L2), and “dorky” to use HMD on streets. In addition, the aesthetics and form factor of the MR devices were also a point of concern, with one comment comparing the experience to strapping “a heavy, awkward thing to the front of my face” (L5). These responses reflect explicit social disapproval of public immersive MR games. They underscore the challenges of overcoming social awkwardness and perceived impropriety. It is crucial to address these concerns to foster widespread adoption and acceptance of this emerging form of entertainment in public spaces.
@@ 5.2.4. Safety and security concerns (7/63)
Many people expressed practical concerns about the potential safety issues that immersive MR street play can result in (5/63). There is a common concern about collision accidents due to the players being too immersed in the MR game and ignoring the real-world context among online respondents. One person commented about the MOFA videos, “I do worry that one of my kids is going to get trampled by an adult who can’t see because he’s too busy slaying a dragon” (L17). Another major concern is about the HMD device being stolen (2/63). People worried that putting an iPhone in the headset would bring a “significant risk of phone theft in major cities” (L24).
@@ 5.3.4. Frustration caused by technological inadequacies (12/65)
Participants also complained about the lack of prompt feedback to their body movement. “It didn’t look like I sent it out with my own hands. The angle of my hand, the angle of my body, and the angle of my head will not affect the result” (P7). Similarly, another participant (P8) stated, “I have no way of judging how precise I am going to be, it’s hard to aim.

### v3

@@ 1. Introduction
Major technology companies such as Apple and Meta are rapidly advancing see-through Mixed Reality Head-Mounted Display (MRHMD) technology: Apple’s Vision Pro (Apple Inc., 2023) offers video see-through mixed reality in a headset form factor, while Meta’s Orion (Meta Platforms, Inc., 2024) boasts optical see-through capabilities in an eyeglasses‑style form factor. Meta CEO Mark Zuckerberg has even predicted that “smart glasses will replace phones by 2030” (Birch, 2024). Some scholars forecast that augmented reality could become virtually indistinguishable from physical reality in certain contexts (Nijholt, 2023; Rauschnabel, 2021), underscoring the importance of understanding user experiences and social dynamics surrounding these emerging devices. As Itoh et al. (Itoh et al., 2021) demonstrate, passthrough technologies—whether video see-through or optical see-through—enable users to maintain awareness of their real-world surroundings while moving freely, extending potential usage from controlled private settings to spontaneous public scenarios. This shift raises pressing questions about how ubiquitous MRHMD adoption could transform everyday public life in the near future.
@@ 2.1. Technology-mediated Street Play
Informed by such studies, HCI has developed conceptual frameworks for thinking about designing for interaction in public, including research in the wild (Chamberlain et al., 2012), performance-led research (Benford et al., 2013), and design considerations of bystanders and spectators in public interactions (Reeves et al., 2005; Reeves, 2011). These approaches have been instrumental in understanding and shaping user experiences in outdoor settings. Meanwhile, HCI research has documented numerous social phenomena that arise when people interact with technology in public contexts, such as public displays (O’Hara et al., 2008) and public installations (Balestrini et al., 2016). Among these are the honeypot effect (Wouters et al., 2016), where an initial user attracts onlookers and encourages them to engage; showmanship and performance, where users become aware of an audience and amplify their actions  (Benford and Giannachi, 2011; Peltonen et al., 2008; Reeves et al., 2005); social mirroring, in which newcomers imitate observed behaviors (Peltonen et al., 2008). Researchers have also noted social discomfort (Benford et al., 2012), social facilitation (improved performance in front of others) (Peltonen et al., 2008; Miller et al., 2019), social awkwardness (Stefanidi et al., 2025; Deterding, 2018; Huggard et al., 2013; Press and Erel, 2023), and inhibition (anxiety or reluctance to participate publicly) (Miller et al., 2019).
@@ 2.2. Technology-supported Co-located Bodily Play
In recent years, HCI research has undergone a “somatic turn” (Loke and Schiphorst, 2018). This shift places the human body at the center of interactive technology design, moving beyond traditional user-centered approaches (Dix, 2017) to focus on the body as the primary site of interaction and experience (Höök, 2018). Merleau-Ponty argues that the human body is fundamental to perception and action (Merleau-Ponty, 2008), with proprioception and kinesthesia serving as core movement-related senses (Zhou et al., 2021). Homewood et al. (Homewood et al., 2021) outline three evolving conceptions of the body in HCI: the individual body (embodied interaction), multiple bodies (plurality of experiences), and more-than-human bodies (entangled assemblages). These concepts, particularly entangled assemblages (Forlano, 2017), intercorporeality (Meyer et al., 2017; Stepanova and Riecke, 2024; Patibanda et al., 2024), and post-phenomenology (Wakkary et al., 2018), provide theoretical frameworks for exploring technology-supported co-located bodily play.
@@ 2.4. Mixed Reality Concerns in Public Spaces
In recent years, MRHMDs with passthrough functionality (Itoh et al., 2021) have sparked a notable shift from private, controlled settings to more public, spontaneous environments. These devices, including video and optical see-through technologies such as Apple Vision Pro, Magic Leap, Microsoft HoloLens, and Meta Orion, enable users to view their physical surroundings while interacting with virtual elements (Radu et al., 2021; Bailenson et al., 2024). The rise in MRHMD usage in public settings raises social concerns for both wearers and non-wearers (Gugenheimer et al., 2017; Sapkota et al., 2021). The social acceptance (Koelle et al., 2020) of MRHMD usage in public spaces tends to decline significantly (Koutromanos and Kazakou, 2023; Profita et al., 2016; Sun and Schöning, 2025; Schwind et al., 2018; Eghbali et al., 2019; Gugenheimer et al., 2019, 2022; O’Hagan et al., 2023a; Lebeck et al., 2018) due to three main factors:
@@
MRHMDs grant wearers capabilities—such as data overlays or video recording—that non-wearers cannot see or fully understand. This imbalance can negatively affect social experiences, privacy, and safety (Chung et al., 2023). Denning et al. (Denning et al., 2014) examined bystander privacy in AR, while O’Hagan et al. (O’Hagan et al., 2021, 2023b) studied bystanders’ privacy concerns and awareness regarding AR glasses. A notable example of these social tensions is Google Glass, which faced criticism for prioritizing user capabilities over broader social impacts. This led to the term “Glasshole”, illustrating bystanders’ perception of antisocial behavior by Google Glass wearers (Due, 2015).
@@
When MRHMDs are used “on the go” in public spaces (Schwind et al., 2018), such as urban streets (Guo et al., 2024), public libraries (Kaeder et al., 2024), public transit (Bajorunaite et al., 2025, 2023; Medeiros et al., 2022; Bajorunaite et al., 2024; Medeiros et al., 2024a), and airplanes (Ng et al., 2021; Williamson et al., 2019), social acceptance, trust, security, and privacy become even more pressing (Katins et al., 2024; Millard et al., 2024; Roesner et al., 2014; Zhu et al., 2024b; Abraham et al., 2022). Interactions and movement in public traffic environments can pose additional risks, especially when MRHMD wearers navigate between physical and virtual elements. These interactions can intensify existing power imbalances and increase the potential for harms and social friction with bystanders.
@@ 2.6. Methods for Exploring Future Mixed Reality
Traditional design futuring (Fry, 2018; Kozubaev et al., 2020; Celik and Kaya, 2025; Lindley and Coulton, 2015) practices, including speculative design (Auger, 2013; Dunne and Raby, 2013; Barendregt and Vaage, 2021), design fiction (Sterling, 2009; Blythe, 2014), experiential futures (Candy and Dunagan, 2017), critical design (Bardzell et al., 2012), material speculation (Wakkary et al., 2015), typically employ fictional or low‑fidelity artifacts (Zhu et al., 2024a; Gaver, 2012), such as diegetic prototypes (Kirby, 2010), mock‑ups (Blythe et al., 2016), scripted scenarios (Candy and Dunagan, 2017), games (Coulton et al., 2016; Brown et al., 2016), or concept videos (Briggs et al., 2012; Odom et al., 2024; Sturdee et al., 2016) to provoke reflection, to imagine alternative worlds, and to foreground the discursive power. Such formats, however, often leave audiences on the far side of what Candy calls an “experiential gulf” (Candy, 2010): viewers discuss a hypothetical future rather than live it, producing debate rather than situated behavioral data (Ringfort-Felner et al., 2025). Elsden et al. (Elsden et al., 2017) further caution that lab‑based user enactments (Odom et al., 2012) can break the suspension of disbelief. To address this gap and facilitate experiential participation, Speculative Enactments (Elsden et al., 2017) stage scenarios as lived experiences that foreground real‑world consequences. Building on this, Immersive Speculative Enactments (Simeone et al., 2022) leverage virtual reality technologies to place participants in open‑ended situations that invite authentic action rather than following fixed linear narratives, as is common in Immersive Design Fiction (McVeigh-Schultz et al., 2018). Cools et al. (Cools et al., 2024) offer an Immersive Speculative Enactment example in which participants experience fictional MR contact lens scenarios through a multi-user VR simulation, generating experiential insights into the technology’s potential impact on daily life.
@@ 3.3. System Implementation
We connected all co-located devices using the InstantCopresence toolkit (Hu et al., 2023a). The toolkit enables spontaneous mixed reality sessions when users scan a QR code on the host device. QR code scanning enables device co-location by calculating the relative position of client devices to the host’s QR code, allowing multiple devices to sync their coordinate systems based on their positions relative to the host origin. The network layer uses Multipeer Connectivity—Apple’s peer-to-peer local networking technology that powers AirDrop—to provide local networking with latency under 20 ms. This enables street play anywhere, even in rural areas without cellular or Wi-Fi coverage. An auto-sync system keeps virtual object data consistent across devices, ensuring all participants see the same virtual elements. Spectators can view the augmented content from a third-person perspective using mobile devices like iPads (see Fig. 4).
@@ 4.1.2. Procedure
    Conclude gameplay: Participants could play for their desired duration. However, as the MOFA games involved vigorous physical activity, players typically concluded after 3–5 minutes due to physical fatigue from bodily exertion. When bystanders wished to join, researchers rotated them into the session. This fluid grouping resembled real-world street play, where people can spontaneously join or leave a game.
@@ 4.1.4. Safety and Ethical Considerations
We avoided any hazardous locations. Although privacy is not expected in public spaces, we followed strict ethical guidelines during data collection. Our recordings focused on gameplay, and we avoided recording people unnecessarily. Participants or bystanders could opt out before interviews. No minors were interviewed; a few minors who tried the game did so under parental supervision, but we did not include their feedback in our dataset beyond general observations.
@@ 4.2. Desktop Research of Online Forum Reviews
In addition to on-site data, we wanted to capture broader public perceptions of IMRSP—including those of people who had not played the game but saw it online. To do this, we conducted a desktop analysis of online discussions about our game probe. We posted a short gameplay video of MOFA and a description of the concept on a popular tech-enthusiast forum, inviting feedback and discussion. Over several weeks, this thread garnered considerable attention and commentary. We collected all responses from this forum thread, resulting in 14,919 words of raw text across numerous comments. Fig. 7 shows snippets from the online forum threads. We then filtered this dataset to focus on relevant content, specifically, comments that addressed the social implications, user experience, or concerns about MR street play (as opposed to off-topic remarks about technology or economics). After filtering out unrelated discussion, we had 1,325 words of substantive comments from 31 unique online respondents (L1–L31). These comments ranged from short quips to multi-sentence opinions (5 to 204 words each).
The online respondents had all viewed the same gameplay footage that our real-world participants had experienced (via the posted video). Their perspectives are thus those of observers envisioning the experience, rather than firsthand players. Not being physically present or involved with researchers, their feedback was not influenced by any demand characteristics or desire to please us. This provided us with more candid feedback and skepticism than we typically observe on‑site. Some commenters were enthusiastic about the idea, while others were critical or concerned. We treated this as an additional qualitative dataset that complements the interview data: on-site participants provide insight into actual experience and immediate personal feelings, whereas online responses reveal more reflective or hypothetical viewpoints, including critiques from a social acceptance standpoint.
@@ 5.2.1. Social awkwardness and perceived impropriety (13/63)
Thirteen respondents (13/63), all from the online forum data collection, expressed feelings of awkwardness as their main impression of the game, with some feeling offended by the prospect of engaging in immersive MR street play. These comments questioned the social acceptance level of wearing the device in public, particularly regarding the potential impact on non-participants. For example, “I’m not sure if I’d like to walk around the streets wearing them though—it would likely be incredibly annoying to other people” (L1). L19 criticized the play as “disregard for culture and obvious social media reactions.” One comment expressed empathy for passersby who might be unintentionally captured in the middle of others’ game space, noting that they “don’t look too pleased” (L26). The impropriety of engaging in the game in certain contexts was also highlighted, with L10 commenting, “It seems pretty impolite to be LARPing in a busy train station”. The remaining comments expressed more straightforward emotions towards this phenomenon, accusing the MR street players of being “a nuisance to everyone” (L27), and calling them “complete idiots to play in public places” (L2), and “dorky” to use HMD on streets. In addition, the aesthetics and form factors of the MR devices were also a point of concern, with one comment comparing the experience to strapping “a heavy, awkward thing to the front of my face” (L5). These responses reflect explicit social disapproval of public immersive MR games. They underscore the challenges of overcoming social awkwardness and perceived impropriety. It is crucial to address these concerns to foster widespread adoption and acceptance of this emerging form of entertainment in public spaces.
@@ 5.2.4. Safety and security concerns (7/63)
Many people expressed practical concerns about the potential safety issues that immersive MR street play may result in (5/63). There is a common concern about collision accidents due to the players being too immersed in the MR game and ignoring the real-world context among online respondents. One person commented about the MOFA videos, “I do worry that one of my kids is going to get trampled by an adult who can’t see because he’s too busy slaying a dragon” (L17). Another major concern is about the HMD device being stolen (2/63). People worried that putting an iPhone in the headset would bring a “significant risk of phone theft in major cities” (L24).
@@ 5.3.4. Frustration caused by technological inadequacies (12/65)
Participants also complained about the lack of prompt feedback to their body movements. “It didn’t look like I sent it out with my own hands. The angle of my hand, the angle of my body, and the angle of my head will not affect the result” (P7). Similarly, another participant (P8) stated, “I have no way of judging how precise I am going to be, it’s hard to aim.